Status Report

NASA ISS On-Orbit Status 26 March 2011

By SpaceRef Editor
March 26, 2011
Filed under , , ,
NASA ISS On-Orbit Status 26 March 2011

All ISS systems continue to function nominally, except those noted previously or below. Saturday – Crew off duty.

Upon wake-up, CDR Kondratyev performed the regular daily check of the aerosol filters at the Russian Elektron O2 (oxygen) generator which Maxim Suraev had installed on 10/19/09 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). [Dmitri will inspect the filters again before bedtime tonight, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]

As generally on Saturdays, the crewmembers joined in conducting the regular weekly three-hour task of thorough cleaning of their home, including COL (Columbus Orbital Laboratory) and Kibo JPM (JEM Pressurized Module). [“Uborka”, usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the sleep stations with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the uborka house cleaning, the CDR completed regular maintenance inspection & cleaning of fan screens in the FGB (TsV2) plus Group E fan grilles (VPkhO, FS5, FS6, VP) in the SM (Service Module) and the Russian POTOK-150MK (150 micron) air purification subsystems (UOV) in the SM & FGB, cleaning the pre-filters with a vacuum cleaner with narrow nozzle attachment and later restarting POTOK in automatic mode.

Dmitri Kondratyev also had another 3 hrs reserved for Progress 41P cargo activities, involving freight unloading, transfers & trash loading, guided by an uplinked “choreography script”.

In preparation for AMS-2 (Alpha Magnetic Spectrometer 2) installation on ULF6, FE-6 Cady Coleman swapped the HRDL (High Data Rate Link) port used up to now for the video MPC (Multi-Protocol Converter) by relocating the OCA data cable HRDL connection from loc. O4 to O1.

Afterwards, Coleman performed the regular bi-monthly reboots of the OCA Router & FS SSC (File Server Station Support Computer) laptops (the latter required to mitigate the problem of iPV and OSTPV (Onboard Short-Term Plan Viewer) not opening up procedures due to a memory leak in the “Tomcat” software). After a period of no less than 10 min later, FE-6 also rebooted the two ISS network servers (SERVER1 & 2).

In the Kibo JPM, Cady prepared MELFI-3 (Minus Eighty Laboratory Freezer for ISS 3) by retrieving 16 +4C ice bricks and inserting them in Dewar 3 trays, plus 4 half-Box Modules for upcoming storage needs.

CDR Kondratyev had ~15 min to transfer his latest video recording to SSC (Station Support Computer) laptop for downlink to TsUP-Moscow. [Dmitri’s Roskosmos “Videoblog” project, recorded by him at his convenience, began on 1/8, with frequent downlinks via OCA (not to exceed 300 Mb),]

Dima also completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers, replacement of EDV-SV waste water and EDV-U urine containers and filling EDV-SV, KOV (for Elektron), EDV-ZV & EDV on RP flow regulator.]

As part of his current VolSci (Voluntary Weekend Science), Nespoli supported the ground in remotely-commanded ERB2 (Erasmus Recording Binocular) activities by turning the payload on and checking it out. [As yesterday, ground controllers at COL-CC (Control Center) then performed ERB2 video file transfers to the EDR (European Drawer Rack) and afterwards downlinked the files via HRDL (High Rate Data Link) at ~8 Mb/s.]

With the Lab CDRA (Carbon Dioxide Removal Assembly) deactivated this morning by ground commanding in support of a scheduled Vozdukh CO2 scrubber performance test, Paolo disconnected the CDRA’s LTL (Low Temperature Loop) return line at the AR2 (Atmosphere Revitalization 2) rack.

CDR & FE-5 -6 had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Dmitri at ~5:45am, Paolo at ~10:10am EDT.

At ~4:30am, Paolo powered up the SM’s amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply) and at 4:35am conducted a ham radio session with students at Comprensivo “G.Manzi”, Civitavecchia, Roma, Italy.

At ~11:15 am, Dima supported three Russian PAO TV events, downlinking messages of greetings to (1) Samara Region representatives plus managers & veterans of the TsSKB-PROGRESS State Research & Production Space Center in Kuibyshev (“…We all know that you deserve a huge credit for this anniversary: Two stages of the Vostok rocket-launcher for the first human flight into space were assembled right here in Kuibyshev. Yuri Gagarin’s flight was made possible because of the work done by your colleagues, employees of many enterprises of Samara….”); (2) participants of a Man-Earth-Space Conference in Kaluga (On April 9-10, the leadership of Kaluga region with the K. E. Tsiolkovsky Cosmonautics Academy are hosting the International Science Conference Man-Earth-Space in Kaluga dedicated to 50th anniversary of the first human flight into space. All senior leaders of the Russian Federation, scientists & various space industry specialists are expected to participate in the conference); and (3) participants of a celebratory gathering in Kaluga (On April 11, the administrative board of Kaluga will conduct a Grand Meeting and a Holiday Concert as a part of 50th anniversary festivities dedicated to Y. A. Gagarin’s space flight. Kaluga & regional leadership, space rocket industry scientists, and specialists will take part in this grand meeting.)

The crew worked out with their regular 2-hr physical exercise on the TVIS treadmill (CDR), ARED advanced resistive exerciser (CDR, FE-5, FE-6), and T2/COLBERT advanced treadmill (FE-5, FE-6). [T2 snubber arm inspection is no longer needed after every T2 session but must be done after the last T2 session of the day.]

WRM Update: A new WRM (Water Recovery Management) “cue card” was uplinked to the crew for their reference, updated with their latest CWC water audit. [The new card (27-0041A) lists 132 CWCs (2,519.5 L total) for the five types of water identified on board: 1. technical water (16 CWCs with 546.7 L, for Elektron electrolysis, incl. 241.7 L in 6 bags containing Wautersia bacteria, 134.2 L in 3 clean bags for contingency use, 170.8 L in 4 bags for transfer into EDV-RP containers via US/RSA-B hose, and 3 empty bags; 2. potable water (no CWCs); 3. iodinated water (104 CWCs with 1,868.9 L for reserve, of which 642.2 L in 35 CWCs are listed as “expired”; 4. condensate water (76.6 L in 10 bags incl. 7.1 L in 1 bag to be used only for OGA, plus 5 empty bags); and 5. waste/EMU dump and other (27.3 L in 2 CWCs from hose/pump flush). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.]

Weekly Science Update (Expedition Twenty-Seven — Week 2).

2D NANO Template (JAXA): MELFI insertion only.

3D SPACE: No report.

AgCam (Agricultural Camera): No report.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.


AMS-02 (Alpha Magnetic Spectrometer): AMS was installed in canister for transport to the launch pad.

APEX (Advanced Plant Experiments on Orbit) -Cambium: No report.

APEX-TAGES (Transgenic Arabidopsis Gene Expression System): No report.

Asian Seed 2010 (JAXA): Photo session is on the task list.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): No report.

BIOLAB (ESA): Return of BIOLAB Handling Mechanism (HM) Gripper on 24S, pending confirmation of good reception.

BIORHYTHMS (JAXA, Biological Rhythms): No report.

BISE (CSA, Bodies in the Space Environment): No report.


BXF-NPBX/MABE (Boiling eXperiment Facility-Nucleate Pool Boiling Exp./Microheater Array Boiling Exp.): “Paolo, this week you installed BXF into the MSG Work Volume. Following activation, we completed functional and video checkouts and heater calibrations for both science investigations, MABE and NPBX. Pending the resolution of some remaining issues, which we expect to be resolved with some brief checkouts, we look forward to beginning science test point operations next week. Thank you for your efforts on our behalf.”

CARD (Long Term Microgravity: Model for Investigating Mechanisms of Heart Disease, ESA): “Thank you, Paolo, for performing the CDL HLTA test. This is now ready to be used for the CARD experiment.”

CARDIOCOG-2: Complete.

CB (JAXA Clean Bench): No report.

CBEF (JAXA Cell Biology Experiment Facility)/SPACE SEED: No report.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): No report.

CERISE (JAXA): No report.

CCF (Capillary Channel Flow): No report.

CFS-A (Colored Fungi in Space-A, ESA): No report.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Ongoing.

SpaceRef staff editor.